OpenStudy (anonymous):

If you are solving y 2+2y=48 by completing the square, the next line would be

OpenStudy (anonymous):

What number will make y^2 + 2y + __ a perfect square? Take coefficient of y, divide by 2, then square it = (2/2)^2 = 1 So add 1 to both sides, since y^2 + 2y + 1 = (y + 1)^2 y^2 + 2y + 1 = 49